• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name folts

    Log entry time 18:03:54 on April 26, 2007

    Entry number 200181

    keyword=target collimator

    At Doug’s request we attempted to move the racetrack target collimators into position.

    After the initial remote failed to give the correct indicators we entered the Hall to assess and found that the collimators had moved, but not far enough in to hit the limit switch.

    Switched to local mode and moved the collimators out and noted that they no longer actuated the switches for the out position. (the switches were made prior to initial actuation).

    Cycled in and out half a dozen times with no improvement. Raised the actuation gas pressure from 100 psi to 120 psi with no noticeable effect.

    Called Javier for guidance and verification.

    Javier confirmed that initial assessment was correct.

    Boosted pressure to 140 psi with no noticeable effect.

    In the open position, the collimators appears to be out of the left spectrometer acceptance. The unknown is the effect of parallax since the target must be viewed from bottom to top through the view ports.

    In the out position the stops seem to be hitting the switch but not actuating them.

    In the in position the collimators are at least 1 cm from the correct position.

    Installed camera on pressure gage to verify that there is gas pressure in the actuation cylinders holding the collimators in the open position.

    DO NOT ATTEMPT TO ACTUATE THE TARGET COLLIMATORS

    Due to the control circuit logic, the collimator system has to remain in remote mode of operation in order to maintain pressure to the open side.



    A copy of this log entry has been emailed to: fansler@jlab.org